OpenStudy (anonymous):

Find the equation of the tangent and the normal to the curve y = x4 - 3x3 + 6x + 2 at the point (2, 6)

OpenStudy (loser66):

take y' , calculate y'(2) =m ( if you are right, it should be 2) plug that m into formula y -6 = m ( x -2) that's it.